Circular No. 7096

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

GRB 990123 R. Lachaume and O. Guyon, Institut d’Astrophysique, Paris, report

their locating the optical candidate at the expected position (cf. IAUC 7094, 7095) on two 30-min CCD exposures with the 1.2-m telescope of the Observatoire de Haute Provence, yielding the following magnitudes (from nine USNO A2.0 catalogue stars), with 1σ uncertainties of ± 0.15 mag: Jan. 24.15 UT, B = 18.40; 24.17, R = 19.45.

E. Ofek and E. M. Leibowitz, Tel Aviv University, write: “From four CCD frames of the object taken at the Wise Observatory between Jan. 24.0617 and 24.1162 UT, we have determined an average R magnitude of 19.87 ± 0.2, with calibration based on four USNO A2.0 stars. The decay parameter in the R band over that time interval is −1.9 ± 0.1. On Jan. 24.0625, the color index of the object was B − R = −0.48 ± 0.25.”

D. D. Kelson, Department of Terrestrial Magnetism, Carnegie Institution of Washington; G. D. Illingworth, University of California at Santa Cruz (UCSC); M. Franx, University of Leiden; D. Magee, UCSC; and P. G. van Dokkum, Kapteyn Astronomical Institute, Groningen, report: “Three 10-min spectra were taken of the proposed optical counterpart to GRB 990123 (IAUC 7094) with the Keck II 10-m telescope on Jan. 24.6 UT. The spectra show the following ultraviolet absorption lines: Fe II (234.4, 237.5, and 238.3 nm), Fe II (258.7 and 260.0 nm), Mg II (279.6 and 280.3 nm), and Mg I (285.2 nm). The redshift is z = 1.61. No strong emission lines were detected between 460 and 920 nm.”

SUPERNOVA 1999J Further to the item on IAUC 7087 regarding the variable in the Large

Magellanic Cloud, A. Becker writes that on Jan. 17.672 UT, it appeared to be near maximum brightness at R = 16.8, V = 17. Discussions between Becker and P. Garnavich (Center for Astrophysics) suggest that the spectrum seems to be most consistent with a peculiar type-Ia supernova, similar to SN 1991T. Garnavich finds a redshift of 10 000 ± 2000 km/s. The spec-

trum is available at http://cfa-www.harvard.edu/~peterg/becker.gif.

R CORONAE BOREALIS Visual magnitude estimates: 1998 Dec. 5.24 UT, 7.9 (A. Pereira, Cabo

da Roca, Portugal); 1999 Jan. 5.23, 7.5 (Pereira); 14.17, 8.0 (K. Hornoch, Lelekovice, Czech Rep.); 18.27, 8.5 (Pereira); 21.25, 9.3 (P. Schmeer, Bischmisheim, Germany); 23.25, 9.5 (Pereira); 23.25, 9.7 (Schmeer).

1999 January 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green